If x+y=40, then what is 2x equal to?
Answer by Maths68(1474)

x+y=40
Multiply both sides by 2
2(x+y)=(2)(40)
2x+2y=80
Subtract 2y from both sides
2x+2y-2y=80-2y
2x=80-2y